# Spacetime Spins: Statistical mechanics for error correction with stabilizer circuits

### v1

@@ I Introduction
Quantum error-correction (QEC) enables quantum computers to operate reliably despite the presence of noise [kitaevQuantum1997, shorFaulttolerant1997, bravyiQuantum1998, knillResilient1998, dennisTopological2002, kitaevFaulttolerant2003, aharonovFaultTolerant2008, gottesmanStabilizer1997, steaneError1996, preskillFaultTolerant1998, poulinStabilizer2005]. The simplest QEC setting is that of a quantum memory, robustly storing logical information via repeated rounds of syndrome extraction. Quantum computation then evolves this with fault-tolerant logical operations that limit the accumulation of errors.
@@
Quantum memories exemplify noise-resilient phases of matter [dennisTopological2002, bombinTopological2010, bombinStrong2012, kubicaThreeDimensional2018, chubbStatistical2021, baoMixedstate2023, fanDiagnostics2024, leeMixedState2025, sangMixedState2024, vennCoherentError2023, behrendsStatistical2025, behrendsSurface2025, baoPhases2024, sangStability2025, wichettePartition2025, englishIsing2025, lavasaniStability2025, chengEmergent2025, jingIntrinsic2025, huangRobust2025, liuCoherent2025, xuPhenomenological2025, wangDecoherenceinduced2025, chenNishimori2025, hauserInformation2025, putzFlow2025, zhuNishimoris2023, ecksteinLearning2025]: The probability of logical errors, given a code and noise model, can be directly expressed through partition functions of statistical mechanical (SM) Ising models [dennisTopological2002]. The error threshold—a critical physical error rate, below which increasing the number of physical qubits leads to a lower logical error rate [kitaevQuantum1997, shorFaulttolerant1997, bravyiQuantum1998, knillResilient1998, dennisTopological2002, kitaevFaulttolerant2003, aharonovFaultTolerant2008]—then emerges as a phase transition between ordered (error correcting) and disordered (non-correcting) phases. Identifying these phase boundaries allows one to assess the theoretical, decoder-agnostic, memory limits.
@@
An important emerging frontier concerns how concepts of phases of matter can capture inherently dynamical aspects of fault tolerance, e.g., syndrome-extraction circuits, logical operations for quantum computations, or indeed intrinsically dynamic codes [hastingsDynamically2021, haahBoundaries2022, tangPhases2025, claesDynamic2025, aasenMeasurement2023, ellisonFloquet2023, davydovaQuantum2024, vuStable2024, aitchisonCompeting2025, vodolaFundamental2022, xuError2025, shawLowering2025, sommersDynamically2025, negariSpacetime2024, ecksteinRobust2024, zhuQubit2023, cainFast2025, martonLattice2025, chirameStabilizing2025, chirameStable2025, takouEstimating2025, zhuStructured2024, gottesmanOpportunities2022, derksDynamical2025, sommersDynamically2025, davydovaUniversal2025, setiawanTailoring2025, derksDesigning2024, williamsonDynamical2025, xuFaulttolerant2025, tanggaraStrategic2024, tanggaraSimple2024, fuError2025, kishonyIncreasing2025]. In this paper, we address this question by showing how to construct SM models for stabilizer circuits—with elementary Clifford gates, resets, and measurements, furnishing the most general fabric behind fault-tolerant constructions.
@@
Leveraging stabilizer dynamics is of increasing focus in practical QEC [bombinUnifying2024, derksDynamical2025, davydovaUniversal2025, setiawanTailoring2025, geherErrorcorrected2024, mcewenRelaxing2023, eickbuschDemonstration2025, debroyLUCI2024, strikisQuantum2023, ankerOptimized2025, wolanskiAutomated2025, martielLowoverhead2025, gidneyLess2023, sahayError2025, serra-peraltaDecoding2025, beverlandFault2024, delfosseSpacetime2023, derksDesigning2024]; examples include designing circuits to overcome hardware limitations [mcewenRelaxing2023, eickbuschDemonstration2025], avoid qubit or coupler dropouts [debroyLUCI2024, strikisQuantum2023, ankerOptimized2025, wolanskiAutomated2025], reduce overheads [martielLowoverhead2025], or turn a static threshold-less code into a dynamical code with thresholds [gidneyLess2023]. Despite these advancements, except for approaches specific to certain codes or circuits [vodolaFundamental2022, tangPhases2025, xuError2025, sommersDynamically2025], or cases where time describes repeated syndrome measurement in phenomenological models [dennisTopological2002, chubbStatistical2021], SM mappings have largely focused on static codes. Our SM models for stabilizer circuits fill this important gap.
@@
In the dynamic setting, QEC and quantum computations correspond to $(d+1)$-dimensional [(d+1)(d+1)D] circuits. Our SM models first represent these as related stabilizer subsystem codes in $(d+1)$-spatial dimensions, known as spacetime codes [baconSparse2017, gottesmanOpportunities2022, delfosseSpacetime2023, derksDesigning2024, pesahFaulttolerant2025, fuSubsystem2025]. Considering generic uncorrelated Pauli noise channels, we then construct a classical Hamiltonian whose partition function—by summing over related error configurations— accounts for fundamental error thresholds and the lowest achievable logical error rates; our SM models are thus of the same status for the dynamical settings as traditional SM models have been for quantum memories.
@@
Mapping stabilizer circuits onto SM models opens up new avenues for analysis beyond those seen in the static case. This framework allows us to compare different compilations of the syndrome-measurement circuits of a QEC code, for example. It also enables statistical mechanical analysis of the logical circuits, such as through transversal gates, lattice surgery, or code switching. A circuit-based analysis of decodability introduces spatially-correlated noise processes—such as “hook errors” from CNOTs between ancilla and data qubits [beverlandFault2024, sahayError2025]—absent from abstracted phenomenological representations of QEC syndrome measurements. These processes are naturally captured by our SM models. Moving SM mappings beyond static codes thus provides useful practical features, beyond its key role in a holistic understanding of fault tolerance and how this defines noise-resilient dynamic phases of matter.
@@
A central technical ingredient in our approach is the introduction of modular spin diagrams, a graphical language of elementary building blocks for constructing SM models for stabilizer circuits. While usable with general noise models, this approach is particularly simple for independent circuit-level $X$-$Z$ noise (where independent $X$ and $Z$ errors can occur before or after any stabilizer operation in the circuit), for which the SM Hamiltonians can be systematically simplified by integrating out certain spins. The spin diagrams we introduce have similarities with those used in ZX-calculus [bridgemanHandwaving2017, weteringZXcalculus2020] with Pauli flows [xuFaulttolerant2025, bombinUnifying2024]; a core difference is that our spin diagrams encode the noise channel, ensuring that simplifying the diagrams preserves the SM partition functions. (See also Ref. [ruschCompleteness2025] for a related construction that uses ZX-diagrams to manipulate circuits while preserving the effects of faults, and Ref. [pesahFaulttolerant2025] for an approach with similar simplification processes but in the context of chain complexes.)
@@
The rest of the paper is structured as follows: Section II presents a brief background on QEC and spacetime codes. Section III outlines the spin models and diagrams for stabilizer circuits. We then apply this to several demonstrative examples. First, in Sec. IV, we study the 1D repetition code, including different syndrome extraction circuits and logical CNOT operations. These examples will help convey the workings of our approach in the simplistic settings of 2D random bond Ising models. In Sec. V we then study the toric code and show how our approach can describe and make qualitative predictions about the performance of recently introduced dynamic syndrome extraction schemes [mcewenRelaxing2023, eickbuschDemonstration2025, claesDynamic2025, williamsonDynamical2025] compared with standard syndrome extraction circuits. We conclude and provide future directions in Sec. VI. The appendices contain derivations and details of the numerical methods.
@@ II.1 Error-correcting codes
Error-correcting codes are designed to protect encoded logical quantum information. In this paper we use subsystem codes [bombinTopological2010], which are generalizations of stabilizer codes [kitaevQuantum1997, kitaevFaulttolerant2003, bravyiQuantum1998, gottesmanStabilizer1997, steaneError1996, preskillFaultTolerant1998, shorFaulttolerant1997, poulinStabilizer2005] and arise naturally in the spacetime code formalism [baconSparse2017, gottesmanOpportunities2022, delfosseSpacetime2023, fuSubsystem2025, derksDesigning2024, pesahFaulttolerant2025]. Information is encoded in a codespace $\mathfrak{C}$: the simultaneous +1+1-eigenspace of a group $\mathcal{S}$ of commuting Pauli operators, “stabilizers”,
@@ II.2 Spacetime codes
The codes described above protect information in a static quantum memory. Quantum computation, however, requires a quantum state to evolve via the unitaries and measurements of a quantum circuit. The spacetime code formalism allows us to interpret this dynamical state as a static subsystem code—the spacetime code—in one higher spatial dimension [baconSparse2017, gottesmanOpportunities2022, delfosseSpacetime2023, fuSubsystem2025, derksDesigning2024, pesahFaulttolerant2025]. [^2] Although the main results of delfosseSpacetime2023 are for spacetime stabilizer codes, their construction also applies to spacetime subsystem codes, as we shall use here. A decoder of the spacetime code can be used to construct an equivalently-performing decoder for the corresponding circuit. We present here a brief overview of spacetime codes; for more detail we refer the reader to delfosseSpacetime2023.
We assume a Clifford circuit of $N$ physical qubits: at every timestep $t=0,1,\ldots,T$ each qubit is either measured with a Pauli operator or is acted on by a Clifford unitary (which may be the identity). This enforces that all measurements and unitaries at the same timestep have disjoint support and thus commute with each other. In-between each timestep, an error channel, cf. Eq. (4), acts on the physical qubits; we say that errors occur at half-integer times $\tau=0.5,1.5,\ldots,T-0.5$. Although this error model does not explicitly include measurement errors, they occur implicitly whenever an error at time $\tau$ anticommutes with a measurement at $\tau+0.5$; this captures the essential features of circuit-level noise without a separate measurement noise parameter. Independent measurement errors may be incorporated by having an additional spacetime qubit attached to each measurement and subjected to bit-flip errors with probability $p_{\text{meas}}$. [^3] This approach was used in gidneyBenchmarking2022, for example.
@@
The spacetime code is a subsystem code; the gauge group defines equivalences between errors in spacetime that have the same effects on the circuit (syndromes and logical operator outcomes). The gauge group is typically generated by two types of spacetime operators [delfosseSpacetime2023]:
@@ III Spacetime spin models
A powerful property of quantum codes is that the probabilities of error cosets can be related to statistical mechanics; models that delineate the limits up to which QEC can work and encapsulate fundamental notions of noise-resilient phases of matter. Initially introduced by dennisTopological2002 for uncorrelated Pauli errors in the toric code, such SM mappings have since encompassed, for example, more stabilizer codes [bombinStrong2012, chubbStatistical2021, kubicaThreeDimensional2018], subsystem codes [bombinTopological2010], coherent errors [vennCoherentError2023, behrendsStatistical2025, behrendsSurface2025, leeMixedState2025, baoPhases2024], and post-selection [englishThresholds2024, englishIsing2025]. We now describe this mapping for general Clifford circuits—extending beyond static codes—subject to noise channels of the form in Eq. (4). In Sec. III.3, we show how to represent these SM models as spin diagrams that can be constructed modularly for each circuit element. These diagrams are particularly simple when restricted to independent $X$-$Z$ noise channels. Section III.4 formalizes the process of simplifying diagrams by removing low-connectivity spins. For a more detailed derivation of this formulation, see Appendix A.2.
We wish to express the probability of an error coset, $\mathbb{P}(\overline{E})$, as a partition function, $\mathcal{Z}_{E}$. To enumerate all errors in the coset $E\mathcal{G}$, we multiply the representative error $E$ by all possible gauge group elements $g\in\mathcal{G}$. We relate this to the SM model by assigning an Ising spin $\sigma_{k}\in\{-1,1\}$ to each gauge generator $g_{k}$; we include $g_{k}$ in $g$ if $\sigma_{k}=-1$. The spin configuration $\bm{\sigma}=[\sigma_{1},\sigma_{2},\ldots,\sigma_{m}]$ thus specifies a group element
@@
encode the reference string $E$ and will create quenched disorder. An important property is that for any error $\epsilon=Eg(\bm{\sigma})$ in the coset, the product $\eta_{i,\tau}(\alpha)\prod_{k}\sigma_{k}$, over all $k$ where $\left\llbracket{[\alpha]_{i,\tau}},{g_{k}}\right\rrbracket=-1$, equals +1+1 if $\epsilon$ commutes with $\alpha]_{(i,\tau)}$, and 1-1 if it anticommutes. We use this product to assign $\epsilon$ the correct probability via dimensionless interaction strengths $K_{i,\tau}(\alpha)$ that weight each product such that $e^{-H_{E}}$ equals $\mathbb{P}(\epsilon)$ (with $H_{E}$ a dimensionless Hamiltonian). [^4] For simplicity, we work with dimensionless quantities. One can equivalently define $J_{i,\tau}(\alpha)=K_{i,\tau}(\alpha)/k_{\text{B}}T$, where $T$ is a temperature to instead use a Hamiltonian in units of energy. Their value is specified by the noise channel, cf. Eq. (4), via a Nishimori condition (see Appendix A.2),
@@
where $\mathcal{P}=\{I,X,Y,Z\}$. [^5] When some $\mathbb{P}(\epsilon)=0$, this equation is not defined; however, by by taking appropriate limits and using the convention $\exp(\ln 0)=0$, the formalism remains well-defined [chubbStatistical2021]. Combining these components together, we have
@@ III.1 Decoding
Since $\mathcal{Z}_{E}=\mathbb{P}(\overline{E})$ for each logical equivalence class, maximum-likelihood decoding, Eq. (8), reduces to finding the coset with maximum partition function—equivalently, minimum free energy $F_{E}=-\ln\mathcal{Z}_{E}$—consistent with the syndrome [bravyiEfficient2014]. The success probability of a decoder, Eq. (6), can be expressed as a (softmax) function of these free energies, [^6] There is not a unique basis for $\mathcal{G}$ and so the Hamiltonian is basis-dependent; the free energy and probability ratios, however, are independent of the basis.
@@
Below threshold, the free-energy diverges in the thermodynamic limit for all $C_{\textsf{s}}L$ ($L\neq I$, and $C_{\textsf{s}}$ the ML Pauli correction) as the ML success probability goes to $1$ [chubbStatistical2021]. Ordered phases of matter are typically characterized by long-range correlations and extensive free-energy costs associated with introducing macroscopic domain walls; the error threshold generally coincides with ordered-disordered phase transitions in the SM model [dennisTopological2002, chubbStatistical2021].
@@
For example, in a stability experiment [gidneyStability2022], one constructs a syndrome-extraction circuit and tests the ability for a decoder to predict the parity of a product of stabilizers after being subjected to errors. A decoder will certainly fail if an error occurs that flips this parity but does not alter the syndrome; one such error is a timelike string of errors that extends across the whole duration of the experiment (cf. Fig. 4b). These undetected timelike errors are coset representatives of some of the spacetime code’s logical equivalence classes. In stability experiments, one typically uses spatial boundary conditions such that no logical qubit is encoded: the code’s $\bar{X}$ and $\bar{Z}$ anticommute with stabilizers on the boundary and do not form logical equivalence classes. The decoder now chooses between only the trivial logical equivalence class $\mathcal{G}$ and those with undetected timelike errors. In contrast, in a memory experiment, one tests the ability for a decoder to preserve the state of an encoded logical qubit by asking it to predict the final eigenvalue of a logical operator after the circuit and comparing it with the true value. Undetected timelike errors add logical equivalence classes and increase the complexity of decoding; one therefore typically uses temporal boundary conditions with spacetime stabilizers that anticommute with these spanning timelike errors to focus on the effect of the static code’s logical equivalence classes. Memory and stability experiments are discussed further in Sec. IV.1, using the repetition code as a guiding example.
@@
Figure 1: The gauge generators (black labels) introduced by various Clifford gates and corresponding spin model ingredients. Each generator is associated with a spin (white square); interactions (colored lines) connect spins according to Eq. (22). Red and blue lines are interactions on $H_{X}$ and $H_{Z}$ respectively. Gridlines show the spacetime locations $(i,\tau)$, with time going up the page. A qubit $i$ idling at integer time $t$, introduces two gauge generators: $X]_{i,t-0.5}[X]_{i,t+0.5}$ and $Z]_{i,t-0.5}[Z]_{i,t+0.5}$. A CNOT has four gauge generators; $g_{1}=[X]_{i,t-0.5}[XX]_{\{i,j\},t+0.5}$ acts on three spacetime qubits and hence its spin has three legs (interactions). A single-qubit measurement $M_{Z}$ (a single-qubit reset, $R_{Z}$, produces the same spins and interactions) has gauge generators with only one interaction in $H_{Z}$ so we draw its spins atop the corresponding spacetime qubit. $M_{Z}$ introduces no spins to $H_{X}$ (and similarly for $M_{X}$ and $H_{Z}$). Two-qubit measurements $M_{ZZ}$ introduce a spin with four interactions and three spins with two-interactions. Shown here is one basis choice; $ZZ]_{\{i,j\},t-0.5}$ could also have been used instead of $ZZ]_{\{i,j\},t+0.5}$.
@@
Figure 2: (Top) An example circuit for a two-qubit ZZZZ measurement via an ancilla qubit, with associated spin diagrams on the right, using the conventions in Fig. 1. Endcaps mark the corresponding termination of the interaction lines (open lines may join up with interactions at earlier/later times.) Errors occur in-between circuit operations; an example error realization is shown. These flip the sign of bonds on the spin model (dashed lines). $Y$ errors flip signs on both $H_{X}$ and $H_{Z}$. (Bottom) Spins with one interaction are integrated out and the interaction removed from $H_{X}$ (marked by crosses). Spins with two remaining interactions are integrated out and their interactions grouped together. The combined interactions are weighted by the number of spacetime locations grouped into that interaction (unlabeled lines are weight one).

### v2

@@ I Introduction
Quantum error-correction (QEC) enables quantum computers to operate reliably despite the presence of noise [70, 104, 21, 72, 37, 71, 3, 56, 107, 93, 92]. The simplest QEC setting is that of a quantum memory, robustly storing logical information via repeated rounds of syndrome extraction. Quantum computation then evolves this with fault-tolerant logical operations that limit the accumulation of errors.
@@
Quantum memories exemplify noise-resilient phases of matter [37, 18, 17, 73, 30, 9, 47, 75, 100, 115, 10, 11, 8, 99, 121, 45, 74, 27, 66, 65, 79, 124, 120, 25, 62, 94, 130, 42]: The probability of logical errors, given a code and noise model, can be directly expressed through partition functions of statistical mechanical (SM) Ising models [37]. The error threshold—a critical physical error rate, below which increasing the number of physical qubits leads to a lower logical error rate [70, 104, 21, 72, 37, 71, 3]—then emerges as a phase transition between ordered (error correcting) and disordered (non-correcting) phases. Identifying these phase boundaries allows one to assess the theoretical, decoder-agnostic, memory limits.
@@
An important emerging frontier concerns how concepts of phases of matter can capture inherently dynamical aspects of fault tolerance, e.g., syndrome-extraction circuits, logical operations for quantum computations, or indeed intrinsically dynamic codes [61, 60, 111, 31, 2, 44, 33, 118, 4, 116, 126, 103, 106, 89, 41, 131, 24, 83, 29, 28, 110, 129, 57, 39, 106, 32, 102, 38, 127, 19, 122, 125, 113, 112, 49, 69]. In this paper, we address this question by showing how to construct SM models for stabilizer circuits—with elementary Clifford gates, resets, and measurements, furnishing the most general fabric behind fault-tolerant constructions.
@@
Leveraging stabilizer dynamics is of increasing focus in practical QEC [20, 39, 32, 102, 51, 84, 43, 35, 108, 6, 123, 82, 52, 98, 101, 15, 36, 38, 127, 19]; examples include designing circuits to overcome hardware limitations [84, 43], avoid qubit or coupler dropouts [35, 108, 6, 123], reduce overheads [82], or turn a static threshold-less code into a dynamical code with thresholds [52]. Despite these advancements, except for approaches specific to certain codes or circuits [116, 111, 126, 106], or cases where time describes repeated syndrome measurement in phenomenological models [37, 30], SM mappings have largely focused on static codes. Our SM models for stabilizer circuits fill this important gap.
@@
In the dynamic setting, QEC and quantum computations correspond to $(d+1)$-dimensional [(d+1)(d+1)D] circuits. Our SM models first represent these as related stabilizer subsystem codes in $(d+1)$-spatial dimensions, known as spacetime codes [7, 57, 36, 38, 91, 50]. Considering generic uncorrelated Pauli noise channels, we then construct a classical Hamiltonian whose partition function—by summing over related error configurations— accounts for fundamental error thresholds and the lowest achievable logical error rates; our SM models are thus of the same status for the dynamical settings as traditional SM models have been for quantum memories.
@@
Mapping stabilizer circuits onto SM models opens up new avenues for analysis beyond those seen in the static case. This framework allows us to compare different compilations of the syndrome-measurement circuits of a QEC code, for example. It also enables statistical mechanical analysis of the logical circuits, such as through transversal gates, lattice surgery, or code switching. A circuit-based analysis of decodability introduces spatially-correlated noise processes—such as “hook errors” from CNOTs between ancilla and data qubits [15, 98]—absent from abstracted phenomenological representations of QEC syndrome measurements. These processes are naturally captured by our SM models. Moving SM mappings beyond static codes thus provides useful practical features, beyond its key role in a holistic understanding of fault tolerance and how this defines noise-resilient dynamic phases of matter.
@@
A central technical ingredient in our approach is the introduction of modular spin diagrams, a graphical language of elementary building blocks for constructing SM models for stabilizer circuits. While usable with general noise models, this approach is particularly simple for independent circuit-level $X$-$Z$ noise (where independent $X$ and $Z$ errors can occur before or after any stabilizer operation in the circuit), for which the SM Hamiltonians can be systematically simplified by integrating out certain spins. The spin diagrams we introduce have similarities with those used in ZX-calculus [23, 114] with Pauli flows [125, 20]; a core difference is that our spin diagrams encode the noise channel, ensuring that simplifying the diagrams preserves the SM partition functions. (See also Ref. [97] for a related construction that uses ZX-diagrams to manipulate circuits while preserving the effects of faults, and Ref. [91] for an approach with similar simplification processes but in the context of chain complexes.)
@@
The rest of the paper is structured as follows: Section II presents a brief background on QEC and spacetime codes. Section III outlines the spin models and diagrams for stabilizer circuits. We then apply this to several demonstrative examples. First, in Sec. IV, we study the 1D repetition code, including different syndrome extraction circuits and logical CNOT operations. These examples will help convey the workings of our approach in the simplistic settings of 2D random bond Ising models. In Sec. V we then study the toric code and show how our approach can describe and make qualitative predictions about the performance of recently introduced dynamic syndrome extraction schemes [84, 43, 31, 122] compared with standard syndrome extraction circuits. We conclude and provide future directions in Sec. VI. The appendices contain derivations and details of the numerical methods.
@@ II.1 Error-correcting codes
Error-correcting codes are designed to protect encoded logical quantum information. In this paper we use subsystem codes [18], which are generalizations of stabilizer codes [70, 71, 21, 56, 107, 93, 104, 92] and arise naturally in the spacetime code formalism [7, 57, 36, 50, 38, 91]. Information is encoded in a codespace $\mathfrak{C}$: the simultaneous +1+1-eigenspace of a group $\mathcal{S}$ of commuting Pauli operators, “stabilizers”,
@@ II.2 Spacetime codes
The codes described above protect information in a static quantum memory. Quantum computation, however, requires a quantum state to evolve via the unitaries and measurements of a quantum circuit. The spacetime code formalism allows us to interpret this dynamical state as a static subsystem code—the spacetime code—in one higher spatial dimension [7, 57, 36, 50, 38, 91]. [^2] Although the main results of Delfosse and Paetznick [36] are for spacetime stabilizer codes, their construction also applies to spacetime subsystem codes, as we shall use here. A decoder of the spacetime code can be used to construct an equivalent decoder for the corresponding circuit. We present here a brief overview of spacetime codes; for more detail we refer the reader to Delfosse and Paetznick [36].
We assume a Clifford circuit of $N$ physical qubits: at every timestep $t=0,1,\ldots,T$ each qubit is either measured with a Pauli operator or is acted on by a Clifford unitary (which may be the identity). This enforces that all measurements and unitaries at the same timestep have disjoint support and thus commute with each other. In-between each timestep, an error channel, cf. Eq. (4), acts on the physical qubits; we say that errors (sometimes called faults in the context of noisy circuits [15]) occur at half-integer times $\tau=0.5,1.5,\ldots,T-0.5$. Although this error model does not explicitly include measurement errors, they occur implicitly whenever an error at time $\tau$ anticommutes with a measurement at $\tau+0.5$; this captures the essential features of circuit-level noise without a separate measurement noise parameter. Independent measurement errors may be incorporated by having an additional spacetime qubit attached to each measurement and subjected to bit-flip errors with probability $p_{\text{meas}}$. [^3] This approach was used in Gidney et al. [53], for example.
@@
The spacetime code is a subsystem code; the gauge group defines equivalences between errors in spacetime that have the same effects on the circuit (syndromes and logical operator outcomes). The gauge group is typically generated by two types of spacetime operators [36]:
@@ III Spacetime spin models
A powerful property of quantum codes is that the probabilities of error cosets can be related to statistical mechanics; models that delineate the limits up to which QEC can work and encapsulate fundamental notions of noise-resilient phases of matter. Initially introduced by Dennis et al. [37] for uncorrelated Pauli errors in the toric code, such SM mappings have since encompassed, for example, more stabilizer codes [17, 30, 73], subsystem codes [18], coherent errors [115, 10, 11, 75, 8], and post-selection [46, 45]. We now describe this mapping for general Clifford circuits—extending beyond static codes—subject to noise channels of the form in Eq. (4). In Sec. III.3, we show how to represent these SM models as spin diagrams that can be constructed modularly for each circuit element. These diagrams are particularly simple when restricted to independent $X$-$Z$ noise channels. Section III.4 formalizes the process of simplifying diagrams by removing low-connectivity spins. For a more detailed derivation of this formulation, see Appendix A.2.
We wish to express the probability of an error coset, $\mathbb{P}(\overline{E})$, as a partition function, $\mathcal{Z}_{E}$, where $E\in\mathcal{P}_{\text{st}}$. To enumerate all errors in the coset $E\mathcal{G}$, we multiply the representative error $E$ by all possible gauge group elements $g\in\mathcal{G}$. We relate this to the SM model by assigning an Ising spin $\sigma_{k}\in\{-1,1\}$ to each gauge generator $g_{k}$; we include $g_{k}$ in $g$ if $\sigma_{k}=-1$. The spin configuration $\bm{\sigma}=[\sigma_{1},\sigma_{2},\ldots,\sigma_{m}]$ thus specifies a group element
@@
encode $E$ and create quenched disorder. An important property is that for any error $\epsilon=Eg(\bm{\sigma})$ in the coset, the product $\eta_{i,\tau}(\alpha)\prod_{k}\sigma_{k}$, over all $k$ where $\left\llbracket{[\alpha]_{i,\tau}},{g_{k}}\right\rrbracket=-1$, equals +1+1 if $\epsilon$ commutes with $\alpha]_{(i,\tau)}$, and 1-1 if it anticommutes. We use this product to assign $\epsilon$ the correct probability via dimensionless interaction strengths $K_{i,\tau}(\alpha)$ that weight each product such that $e^{-H_{E}}$ equals $\mathbb{P}(\epsilon)$ (with $H_{E}$ a dimensionless Hamiltonian). [^4] For simplicity, we work with dimensionless quantities. One can equivalently define $J_{i,\tau}(\alpha)=K_{i,\tau}(\alpha)/k_{\text{B}}T$, where $T$ is a temperature to instead use a Hamiltonian in units of energy. Their value is specified by the noise channel, cf. Eq. (4), via a Nishimori condition (see Appendix A.2),
@@
where $\mathcal{P}=\{I,X,Y,Z\}$. [^5] When some $\mathbb{P}(\epsilon)=0$, the expression for $K_{i,\tau}(\alpha)$ is not defined; however, by taking appropriate limits and using the convention $\exp(\ln 0)=0$, the formalism remains well-defined [30]. Combining these components together, we have
@@ III.1 Decoding
Since $\mathcal{Z}_{E}=\mathbb{P}(\overline{E})$ for each logical equivalence class, maximum-likelihood decoding, Eq. (8), reduces to finding the coset with maximum partition function—equivalently, minimum free energy $F_{E}=-\ln\mathcal{Z}_{E}$—consistent with the syndrome [22]. The success probability of a decoder, Eq. (6), can be expressed as a (softmax) function of these free energies, [^6] There is not a unique basis for $\mathcal{G}$ and so the Hamiltonian is basis-dependent; the free energy and probability ratios, however, are independent of the basis.
@@
Below threshold, the free energy diverges in the thermodynamic limit for all $C_{\textsf{s}}L$ ($L\neq I$, and $C_{\textsf{s}}$ the ML Pauli correction) as the ML success probability goes to $1$ [30]. Ordered phases of matter are typically characterized by long-range correlations and extensive free-energy costs associated with introducing macroscopic domain walls; the error threshold generally coincides with ordered-disordered phase transitions in the SM model [37, 30].
@@
For example, in a stability experiment [55], one constructs a syndrome-extraction circuit and tests the ability for a decoder to predict the parity of a product of stabilizers after being subjected to errors. A decoder will certainly fail if an error occurs that flips this parity but does not alter the syndrome; one such error is a timelike string of errors that extends across the whole duration of the experiment (cf. Fig. 5b). These undetected timelike errors are coset representatives of some of the spacetime code’s logical equivalence classes. In stability experiments, one typically uses spatial boundary conditions such that no logical qubit is encoded: the code’s $\bar{X}$ and $\bar{Z}$ anticommute with stabilizers on the boundary and do not form logical equivalence classes. The decoder now chooses between only the trivial logical equivalence class $\mathcal{G}$ and those with undetected timelike errors. In contrast, in a memory experiment, one tests the ability for a decoder to preserve the state of an encoded logical qubit by asking it to predict the final eigenvalue of a logical operator after the circuit and comparing it with the true value. Undetected timelike errors add logical equivalence classes and increase the complexity of decoding; one therefore typically uses temporal boundary conditions with spacetime stabilizers that anticommute with these spanning timelike errors to focus on the effect of the static code’s logical equivalence classes. Memory and stability experiments are discussed further in Sec. IV.1, using the repetition code as a guiding example.
@@
Figure 1: The gauge generators (red and blue labels) from Clifford gates and their corresponding spin model ingredients. Each generator is associated with a spin (white square); interactions (colored lines) connect spins according to Eq. (22). Red and blue lines are interactions on $H_{X}$ and $H_{Z}$ respectively. Gridlines show the spacetime locations $(i,\tau)$, with time going up the page. A qubit $i$ idling at integer time $t$ introduces two gauge generators: $X]_{i,t-0.5}[X]_{i,t+0.5}$ and $Z]_{i,t-0.5}[Z]_{i,t+0.5}$. A CNOT has four gauge generators; $g_{1}=[X]_{i,t-0.5}[XX]_{\{i,j\},t+0.5}$ acts on three spacetime qubits and hence its spin has three legs (interactions). A single-qubit measurement $M_{Z}$ (a single-qubit reset, $R_{Z}$, produces the same spins and interactions) has gauge generators with only one interaction in $H_{Z}$ so we draw its spins atop the corresponding spacetime qubit. $M_{Z}$ introduces no spins to $H_{X}$ (and similarly for $M_{X}$ and $H_{Z}$). Two-qubit measurements $M_{ZZ}$ introduce a spin with four interactions and three spins with two-interactions. Shown here is one basis choice; $ZZ]_{\{i,j\},t-0.5}$ could also have been used instead of $ZZ]_{\{i,j\},t+0.5}$.
@@
Figure 2: (Top) An example circuit for a two-qubit ZZZZ measurement via an ancilla qubit, with associated spin diagrams on the right, using the conventions in Fig. 1. Endcaps mark the corresponding termination of the interaction lines (open lines may join up with interactions at earlier/later times.) Errors occur in-between circuit operations; an example error realization is shown. These flip the sign of bonds on the spin model (dashed lines). $Y$ errors flip signs on both $H_{X}$ and $H_{Z}$. (Bottom) The simplified spin diagram derived by applying the two rules in Fig. 3.
Figure 3: Two rules used to simplify spin diagrams (and their associated Hamiltonians). (1) Spins with one interaction are integrated out and the interaction removed from $H_{\alpha}$ (marked by crosses). (2) Spins with two interactions are integrated out and their interactions grouped together. The combined interaction is weighted by the number of spacetime locations grouped together (unlabeled lines are weight one).